A father is now four times as old as his son.If the sum of their ages ten years ago was sixty.find their present ages ?
Answer by nerdybill(7384)

A father is now four times as old as his son.If the sum of their ages ten years ago was sixty.find their present ages ?
.
Let x = son's age
then
4x = father's age
.
From:"the sum of their ages ten years ago was sixty" we get:
(x-10) + (4x-10) = 60
x - 10 + 4x - 10 = 60
5x - 20 = 60
5x = 80
x = 16 years (son's age)
.
father's age:
4x = 4(16) = 64 years (father's age)